Use the given degree of confidence and sample data to construct a confidence interval for the population proportion p. Round to three decimal places.

Of 89 adults selected randomly from one​ town, 66 have health insurance. Find a​ 90% confidence interval for the true proportion of all adults in the town who have health insurance.
Answer #1

proportion p = 66/89

standard error = sqrt( p *(1-p)/n)

margin of error = z-score * standard error

= 1.645 * sqrt( (66/89) * (23/89)/89)

= 0.07633

lower limit = (66/89) - 0.07633 = 0.6652

upper limit = (66/89) + 0.07633 = 0.8179